# On the reconstruction map in JT gravity

### v1

@@ 1 Introduction
Recently, a precise reconstruction map [1] was proposed in Jackiw-Teitelboim (JT) gravity [2, 3]. Our goal is to review this reconstruction map, describe some of its features – both desirable and (from our perspective) perhaps undesirable – and then propose a modified version without the ‘undesirable’ features. We then use this new reconstruction map to learn things about non-perturbative quantum gravity and the dynamics of quantum wormholes on long time scales, using techniques from quantum ergodicity theory [4].
@@ 1.1 Summary of previous work
JT gravity [2, 3] is a 1+1 dimensional theory of gravity defined by the action
@@
In this number of dimensions, the first term is topological, and a larger parameter $S_{0}$ serves to suppress contributions from non-trivial spacetime topologies to a path integral computation. Consider the theory at $S_{0}\to\infty$, which we will call $\mathrm{JT}_{\infty}$. All solutions in this theory have the same topology of two asymptotic boundaries connected by a wormhole behind the horizon of two black holes. The only dynamical degree of freedom is the length of this wormhole [5]. Quantized $\mathrm{JT}_{\infty}$ is equivalent to a particle moving in one dimension with an exponential potential: it has a Hilbert space $\mathcal{H}_{0}$ and Hamiltonian $H_{0}$ given by
where $x$ is the wormhole length [5]. This Hamiltonian has eigenstates |$\ket{E}_{0}$ for $E>0$, with wavefunctions
$\phi_{E}(x):=\braket{x}{E}_{0}=4\mathrm{K}_{\mathrm{i}\sqrt{8E}}(4\mathrm{e}^{-x/2})$ (1.3)
@@
$\braket{E}{E^{\prime}}_{0}=\frac{\delta(E-E^{\prime})}{\rho_{0}(E)}~{},~{}~{}~{}~{}\int\limits_{0}^{\infty}\mathrm{d}E\;\rho_{0}(E)\ket{E}\bra{E}_{0}=\mathbb{1}_{0}~{},$ (1.4)
@@
At finite $S_{0}$, the topology can fluctuate. When computing with the path integral, we must sum over different topologies [6, 7]. This theory has a known dual description as a random matrix model [7], where each matrix can be regarded as a Hamiltonian in a Hilbert space $\mathcal{H}$, with discrete spectrum $E_{1},E_{2},...$ with typical spacing $\Delta E\simeq\mathrm{e}^{-S_{0}}$ and leading order density of states $\mathrm{e}^{S_{0}}\rho_{0}(E)$. The gravitational path integral calculates averaged quantities over this ensemble of random matrix theories.
Now consider a single draw $H$ from this ensemble. We will follow [1] and consider $\mathrm{JT}_{\infty}$ as an effective description of that fundamental theory. We can then ask how the predictions of $\mathrm{JT}_{\infty}$ are corrected in the fundamental theory. In particular, let us consider the wormhole length starting in the Hartle-Hawking state
@@
where $Z(\beta)$ is the partition function. $\mathrm{JT}_{\infty}$ predicts $\braket{x(t)}$ grows linearly forever in $t$. What does the fundamental theory predict? This question has a long history [8, 9, 10, 1], and some general comments can be made. First, we should expect that the true length matches the prediction of the effective theory for a while – any short experiment shouldn’t be able to resolve the energy differences. However, eternal growth can’t happen for any reasonable choice of $\hat{x}$: in the fundamental theory, $|\beta\rangle$ has non-negligible support on a finite $\mathrm{e}^{S}$ number of energy eigenstates, where $S$ is the thermal entropy of the state. Assuming the matrix elements $\langle E_{n}|\hat{x}|E_{m}\rangle$ are finite, the length must eventually stop growing. In fact, it is expected to plateau (rather than, say, to coherently decrease in length back to its starting size), with small fluctuations around the plateau value for extremely long times after the Heisenberg time, $t_{\mathrm{H}}\sim\mathrm{e}^{S}$.
That said, it is one thing to know $\mathrm{JT}_{\infty}$ is wrong at late times, and another thing to actually compute the finite $S_{0}$ corrections. Explicit computation of these non-perturbative corrections was a major goal of [1]. The first hurdle they encountered is to find an answer to the following question:
Given a $\mathrm{JT}_{\infty}$ state |$\ket{\psi}_{0}$ and operator $\hat{x}_{0}$, what are the corresponding state and operator in the fundamental theory?
The starting point of [1] is to define a holographic map [^1] The presentation of these ideas in [1] is slightly different, but the physics is equivalent.
@@
$\braket{x}{x^{\prime}}_{0}=\delta(x-x^{\prime})~{}.$ (1.8)
At finite $S_{0}$, [1] argues that the inner product is modified to
$\braket{x}{x^{\prime}}=\mathrm{e}^{-S_{0}}\sum_{i=1}^{\infty}\phi_{E_{i}}(x)\phi_{E_{i}}(x^{\prime})~{}.$ (1.9)
@@
$\braket{x}{x^{\prime}}_{0}=\int\mathrm{d}E\;\rho_{0}(E)\phi_{E}(x)\phi_{E}(x^{\prime})\longrightarrow\mathrm{e}^{-S_{0}}\sum_{i=1}^{\infty}\phi_{E_{i}}(x)\phi_{E_{i}}(x^{\prime})=:\braket{x}{x^{\prime}}~{}.$ (1.10)
@@
[…]
Technically, this $V$ is still ambiguous under $V\to UV$ for any unitary $U$. We follow [1] and fix this ambiguity by choosing $V$ to be the one satisfying both (1.11) and which is equivariant under time evolution:
@@
Any reconstruction map we define will require extra input. In [1], it was proposed that we should define
@@
This map has some nice properties. One is just that it is a natural $R^{*}$ to write down, if all that’s known is $V$. Moreover, the authors of [1] used this $R^{*}$ to reconstruct a version of the wormhole length operator, and found the desired qualitative behavior: linear growth for a time $\mathrm{e}^{S}$ followed by a plateau.
However, this map (1.16) also has some undesirable properties (in our view). One is that it does not reconstruct simple operators in a nice way. In particular, there are normalization ambiguities that arise even with the identity operator, which we would like to avoid. Consider the reconstruction of the identity $\mathbb{1}_{0}=\int_{-\infty}^{\infty}\mathrm{d}x\,\ket{x}_{0}\bra{x}_{0}$
$R^{*}(\mathbb{1}_{0})=\int\limits_{-\infty}^{\infty}\mathrm{d}x\,\ket{x}\bra{x},$ (1.17)
where we write |$\ket{x}:=V\ket{x}_{0}$. But then
$\braket{x}{R^{*}(\mathbb{1}_{0})}{x}=\int\limits_{-\infty}^{\infty}\mathrm{d}x^{\prime}\;\lvert\braket{x}{x^{\prime}}\rvert^{2}=\infty~{},$ (1.18)
because $\lvert\braket{x}{x^{\prime}}\rvert$ | is generally $O(\mathrm{e}^{-S_{0}})$. This demonstrates $R^{*}(\mathbb{1}_{0})$ is not the properly normalized identity operator because again $\braket{x}{x^{\prime}}\simeq\delta_{x,x^{\prime}}+O(\mathrm{e}^{-S_{0}})$. Similarly, $R^{*}(x)$ is ill defined, as are the reconstructions of many other simple operators. The authors of [1] pointed this out, and proposed that we simply ignore the operators with bad reconstructions, or use regulated versions for the ones like $x$ whose physics is important to us.
Another recent paper [11] describes a regularization for the length operator based on restricting the path integral to a microcanonical window. Our predictions will be similar but inequivalent.
@@ 1.2 Summary of our results
Our point of view is that a good reconstruction map $R^{*}$ should satisfy a certain compatibility condition with the holographic map $V$, which we summarize as follows. [^2] We focus here on linear reconstruction maps, even though in general holographic systems the reconstruction maps can be non-linear because the holographic map is non-isometric [12, 13]. This restriction is acceptable in this setting of pure JT, without matter, as long as we only demand the reconstructions work on a restricted set of states that we call semiclassical. We return to JT with matter in Section 6.3. Let $S$ be some subset of states in $\mathcal{H}_{0}$, which are morally the “good, semiclassical” states that the effective description should be a valid description of. Let $K_{S}$ be the analog for operators, i.e. the subset of $\mathrm{B}(\mathcal{H}_{0})$ that we wish to regard as “good and semiclassical.” Presumably $K_{S}$ should include the operators that preserve $S$, like the identity.
Roughly speaking, we say that $R^{*}$ satisfies the compatibility condition if for a given $V$ and any |$\ket{\psi}\in S$ and $O\in K_{S}$,
$\bra{\psi}V^{\dagger}R^{*}(O)V\ket{\psi}$ $\simeq\bra{\psi}O\ket{\psi}~{}.$ (1.19)
@@
The reconstruction map (1.16) from [1] does not satisfy this compatibility condition. One natural set of semiclassical states $S$ are the Hartle-Hawking state and short time evolutions of it. The inner product between these states is approximately preserved by $V$ (the corrections from higher topologies are very small), so we would expect $R^{*}(\mathbb{1}_{0})\simeq\mathbb{1}$, but as we saw in (1.18) this is not the case.
@@
In Section 4 we use this reconstruction map to study the length of the wormhole at long times. An unexpected analogy between quantum JT gravity and quantum ergodicity theory [4] is useful, and allows us to make new quantitative predictions about non-perturbative quantum gravity. Firstly, we find that the average wormhole length $\langle x(t)\rangle$ is not monotonically increasing in time: at times $t\sim t_{\mathrm{H}}\sim\mathrm{e}^{S(E)}$, $\langle x(t)\rangle$ grows above its late-time value before decreasing back. This is in tantalizing agreement with an independent recent calculation [14], to which we will compare our results in Section 6.2. An explicit numerical calculation of the wormhole dynamics, describing both the quantum wave function and the averaged wormhole length, is depicted in Fig. 1. Secondly, we predict the magnitude of quantum fluctuations in wormhole length, finding that
@@
at short times; here $\Delta x(0)$ is the initial uncertainty in position and $x_{\mathrm{sat}}$ is the late time expectation value of length. This is a clear prediction of our formalism, and a consequence of the random matrix statistics of the fundamental JT Hamiltonian [7]. Thirdly, we predict the late time asymptotics of the time-evolved Hartle-Hawking state (1.6): crudely speaking, the probability that the length is measured to lie in the interval $[x,x+\mathrm{d}x]$ is
@@ 2.1 The line as an effective description
and infinite potential walls at 00 and $L$. Say we are mainly interested in a state supported near $x=L/2$, dying off very quickly relative to $L$, such as (schematically)
@@
Intuitively, this is a good effective description because we can identify the |$\ket{x}$ states and the $\hat{x}$ operators in the two systems, and then the expectation values match (for a while): $\braket{\hat{x}}_{\mathrm{line}}\approx\braket{\hat{x}}_{\mathrm{int}}$. Note that the effective description has a continuous spectrum while that of the fundamental theory is discrete, analogous to the relationship between semiclassical JT and the dual quantum theory.
@@ 2.2 Codifying the relationship with $V$ and $R^{*}$
$\begin{split}R&:\mathcal{H}_{\mathrm{int}}\to\mathcal{H}_{\mathrm{line}}\\ &:\ket{n}_{\mathrm{int}}\to\sqrt{\frac{2}{L}}\int\limits_{0}^{L}\mathrm{d}x\,\sin\left(\frac{n\pi x}{L}\right)\ket{x}_{\mathrm{line}}\end{split}$ (2.7)
@@
|$\ket{s_{E,\varepsilon}}_{\mathrm{line}}=\frac{\sqrt{2\varepsilon}}{(2\pi)^{\frac{1}{4}}}\int_{-\infty}^{\infty}\mathrm{d}x\,\mathrm{e}^{-x^{2}\varepsilon^{2}/4}\sin(\sqrt{2mE}x)\ket{x}_{\mathrm{line}}~{}.$ (2.12)
These are approximately normalized states of approximately definite energy, with both approximations improving quickly as $\varepsilon\to 0$. We could analogously define states |$\ket{c_{E,\varepsilon}}_{\mathrm{line}}$ which replace the sine with cosine, and together these states for $E>0$ would span $\mathcal{H}_{\mathrm{line}}$.
@@
$V_{\varepsilon}=\frac{(2\pi)^{\frac{1}{4}}}{\sqrt{\varepsilon L}}\sum_{n}\ket{n}_{\mathrm{int}}\bra{s_{E_{n},\varepsilon}}_{\mathrm{line}}~{},$ (2.13)
where $E_{n}$ are the eigenvalues of $H_{\mathrm{int}}$. It is straightforward to show this satisfies conditions (2.10) and (2.11) with error $O(\varepsilon)$. Moreover, $V_{\varepsilon}$ preserves the normalization of wave functions supported entirely in $x\in[0,L]$, up to terms that vanish as $\varepsilon\to 0$. We can make $\varepsilon$ as small as we want, so we will drop $\varepsilon$ going forward, imagining that we are working in the $\varepsilon\to 0$ limit. [^3] For some states, $\lim_{\varepsilon\to 0}V_{\varepsilon}\ket{\psi}$ will not converge. We regard those |$\ket{\psi}$ as outside the regime of validity of the effective theory.
With these maps $V$ and $R^{*}$ in hand, we have formalized the sense in which the particle on a line is an effective description of the particle on the interval. Given some operator and state on the line, these maps tell us what we are supposed to compare to on the interval. They also allow us to determine the regime of validity of the effective description. We say it is valid for a given state |$\ket{\psi}\in\mathcal{H}_{\mathrm{line}}$ and observable $O\in\mathrm{L}(\mathcal{H}_{\mathrm{line}})$ if
@@ 2.3 Lessons for JT
Recall that one idea, from [1], is
@@
$R_{I}^{*}(x_{\mathrm{line}})=\sum_{n,n^{\prime}}\ket{n}\bra{n^{\prime}}_{\mathrm{int}}\frac{2}{L}\int\limits_{-\infty}^{\infty}\mathrm{d}x\,x\sin\left(\frac{n\pi}{L}x\right)\sin\left(\frac{n^{\prime}\pi}{L}x\right)\mathrm{e}^{-x^{2}\varepsilon^{2}/4}$ (2.16)
@@
Note a crucial difference between JT and this line/interval example. In JT, we are starting by knowing $V$ and would like to deduce $R^{*}$. This is trickier than the converse, because $V$ can be straightforwardly determined from $R$ via (2.10) and (2.11). We’ll have to figure out the appropriate $R^{*}$ by different means.

### v2

@@ 1 Introduction
Recently, a precise reconstruction map [Iliesiu:2024cnh] was proposed in Jackiw-Teitelboim (JT) gravity [Jackiw:1984je, Teitelboim:1983ux]. Our goal is to review this reconstruction map, describe some of its features – both desirable and (from our perspective) perhaps undesirable – and then propose a modified version without the ‘undesirable’ features. We then use this new reconstruction map to learn things about non-perturbative quantum gravity and the dynamics of quantum wormholes on long time scales, using techniques from quantum ergodicity theory [dynamicalqergodicity].
@@ 1.1 Summary of previous work
JT gravity [Jackiw:1984je, Teitelboim:1983ux] is a 1+1 dimensional theory of gravity defined by the action
$S_{\mathrm{JT}}=S_{0}\int\mathrm{d}^{2}x\sqrt{-g}R+\int\mathrm{d}^{2}x\sqrt{-g}\Phi(R+2)+\text{(boundary terms)}$ (1.1)
In this number of dimensions, the first term is topological, and a larger parameter $S_{0}$ serves to suppress contributions from non-trivial spacetime topologies to a path integral computation. Consider the theory at $S_{0}\to\infty$, which we will call $\mathrm{JT}_{\infty}$. All solutions in this theory have the same topology of two asymptotic boundaries connected by a wormhole behind the horizon of two black holes. The only dynamical degree of freedom is the length of this wormhole [Harlow:2018tqv]. Quantized $\mathrm{JT}_{\infty}$ is equivalent to a particle moving in one dimension with an exponential potential: it has a Hilbert space $\mathcal{H}_{0}$ and Hamiltonian $H_{0}$ given by
@@
$\phi_{E}(x):=\braket{x|E}_{0}=4\mathrm{K}_{\mathrm{i}\sqrt{8E}}(4\mathrm{e}^{-x/2})$ (1.3)
@@
[…]
@@
At finite $S_{0}$, the topology can fluctuate. When computing with the path integral, we must sum over different topologies [Saad:2018bqo, Saad:2019lba]. This theory has a known dual description as a random matrix model [Saad:2019lba], where each matrix can be regarded as a Hamiltonian in a Hilbert space $\mathcal{H}$, with discrete spectrum $E_{1},E_{2},$... with typical spacing $\Delta E\simeq\mathrm{e}^{-S_{0}}$ and leading order density of states $\mathrm{e}^{S_{0}}\rho_{0}(E)$. The gravitational path integral calculates averaged quantities over this ensemble of random matrix theories.
Now consider a single draw $H$ from this ensemble. We will follow [Iliesiu:2024cnh] and consider $\mathrm{JT}_{\infty}$ as an effective description of that fundamental theory. We can then ask how the predictions of $\mathrm{JT}_{\infty}$ are corrected in the fundamental theory. In particular, let us consider the wormhole length starting in the Hartle-Hawking state
@@
where $Z(\beta)$ is the partition function. $\mathrm{JT}_{\infty}$ predicts $\braket{x(t)}$ grows linearly forever in $t$. What does the fundamental theory predict? This question has a long history [Susskind:2014rva, Iliesiu:2021ari, Stanford:2022fdt, Iliesiu:2024cnh], and some general comments can be made. First, we should expect that the true length matches the prediction of the effective theory for a while – any short experiment shouldn’t be able to resolve the energy differences. However, eternal growth can’t happen for any reasonable choice of $\hat{x}$: in the fundamental theory, $|\beta\rangle$ has non-negligible support on a finite $\mathrm{e}^{S}$ number of energy eigenstates, where $S$ is the thermal entropy of the state. Assuming the matrix elements $\langle E_{n}|\hat{x}|E_{m}\rangle$ are finite, the length must eventually stop growing. In fact, it is expected to plateau (rather than, say, to coherently decrease in length back to its starting size), with small fluctuations around the plateau value for extremely long times after the Heisenberg time, $t_{\mathrm{H}}\sim\mathrm{e}^{S}$.
That said, it is one thing to know $\mathrm{JT}_{\infty}$ is wrong at late times, and another thing to actually compute the finite $S_{0}$ corrections. Explicit computation of these non-perturbative corrections was a major goal of [Iliesiu:2024cnh]. The first hurdle they encountered is to find an answer to the following question:
[…]
The starting point of [Iliesiu:2024cnh] is to define a holographic map [^1] The presentation of these ideas in [Iliesiu:2024cnh] is slightly different, but the physics is equivalent.
@@
$\braket{x|x^{\prime}}_{0}=\delta(x-x^{\prime})~.$ (1.8)
At finite $S_{0}$, [Iliesiu:2024cnh] argues that the inner product is modified to
$\braket{x|x^{\prime}}=\mathrm{e}^{-S_{0}}\sum_{i=1}^{\infty}\phi_{E_{i}}(x)\phi_{E_{i}}(x^{\prime})~.$ (1.9)
@@
$\braket{x|x^{\prime}}_{0}=\int\mathrm{d}E\;\rho_{0}(E)\phi_{E}(x)\phi_{E}(x^{\prime})\longrightarrow\mathrm{e}^{-S_{0}}\sum_{i=1}^{\infty}\phi_{E_{i}}(x)\phi_{E_{i}}(x^{\prime})=:\braket{x|x^{\prime}}~.$ (1.10)
@@
$\braket{x|x^{\prime}}=\braket{x|V^{\dagger}V|x^{\prime}}_{0}~.$ (1.11)
Technically, this $V$ is still ambiguous under $V\to UV$ for any unitary $U$. We follow [Iliesiu:2024cnh] and fix this ambiguity by choosing $V$ to be the one satisfying both (1.11) and which is equivariant under time evolution:
@@
Any reconstruction map we define will require extra input. In [Iliesiu:2024cnh], it was proposed that we should define
@@
This map has some nice properties. One is just that it is a natural $R^{*}$ to write down, if all that’s known is $V$. Moreover, the authors of [Iliesiu:2024cnh] used this $R^{*}$ to reconstruct a version of the wormhole length operator, and found the desired qualitative behavior: linear growth for a time $\mathrm{e}^{S}$ followed by a plateau.
[…]
$|,R^{*}(\mathbb{1}_{0})=\int\limits_{-\infty}^{\infty}\mathrm{d}x\,\ket{x}\bra{x}$, (1.17)
[…]
$\braket{x|R^{*}(\mathbb{1}_{0})|x}=\int\limits_{-\infty}^{\infty}\mathrm{d}x^{\prime}\;\lvert\braket{x|x^{\prime}}\rvert^{2}=\infty~,$ (1.18)
because $\lvert\braket{x|x^{\prime}}\rvert$ is generally $O(\mathrm{e}^{-S_{0}})$. This demonstrates $R^{*}(\mathbb{1}_{0})$ is not the properly normalized identity operator because again $\braket{x|x^{\prime}}\simeq\delta_{x,x^{\prime}}+O(\mathrm{e}^{-S_{0}})$. Similarly, $R^{*}(x)$ is ill defined, as are the reconstructions of many other simple operators. The authors of [Iliesiu:2024cnh] pointed this out, and proposed that we simply ignore the operators with bad reconstructions, or use regulated versions for the ones like $x$ whose physics is important to us.
Another recent paper [Miyaji:2024ity] describes a regularization for the length operator based on restricting the path integral to a microcanonical window. Our predictions will be similar but inequivalent.
@@ 1.2 Summary of our results
Our point of view is that a good reconstruction map $R^{*}$ should satisfy a certain compatibility condition with the holographic map $V$, which we summarize as follows. [^2] We focus here on linear reconstruction maps, even though in general holographic systems the reconstruction maps can be non-linear because the holographic map is non-isometric [Akers:2021fut, Akers:2022qdl]. This restriction is acceptable in this setting of pure JT, without matter, as long as we only demand the reconstructions work on a restricted set of states that we call semiclassical. We return to JT with matter in Section 6.3. Let $S$ be some subset of states in $\mathcal{H}_{0}$, which are morally the “good, semiclassical” states that the effective description should be a valid description of. Let $K_{S}$ be the analog for operators, i.e. the subset of $\mathrm{B}(\mathcal{H}_{0})$ that we wish to regard as “good and semiclassical.” Presumably $K_{S}$ should include the operators that preserve $S$, like the identity.
[…]
@@
The reconstruction map (1.16) from [Iliesiu:2024cnh] does not satisfy this compatibility condition. One natural set of semiclassical states $S$ are the Hartle-Hawking state and short time evolutions of it. The inner product between these states is approximately preserved by $V$ (the corrections from higher topologies are very small), so we would expect $R^{*}(\mathbb{1}_{0})\simeq\mathbb{1}$, but as we saw in (1.18) this is not the case.
@@
In Section 4 we use this reconstruction map to study the length of the wormhole at long times. An unexpected analogy between quantum JT gravity and quantum ergodicity theory [dynamicalqergodicity] is useful, and allows us to make new quantitative predictions about non-perturbative quantum gravity. Firstly, we find that the average wormhole length $\langle x(t)\rangle$ is not monotonically increasing in time: at times $t\sim t_{\mathrm{H}}\sim\mathrm{e}^{S(E)}$, $\langle x(t)\rangle$ grows above its late-time value before decreasing back. This is in tantalizing agreement with an independent recent calculation [Balasubramanian:2024lqk], to which we will compare our results in Section 6.2. An explicit numerical calculation of the wormhole dynamics, describing both the quantum wave function and the averaged wormhole length, is depicted in Fig. 1. Secondly, we predict the magnitude of quantum fluctuations in wormhole length, finding that
@@
at short times; here $\Delta x(0)$ is the initial uncertainty in position and $x_{\mathrm{sat}}$ is the late time expectation value of length. This is a clear prediction of our formalism, and a consequence of the random matrix statistics of the fundamental JT Hamiltonian [Saad:2019lba]. Thirdly, we predict the late time asymptotics of the time-evolved Hartle-Hawking state (1.6): crudely speaking, the probability that the length is measured to lie in the interval $x,x+\mathrm{d}x]$ is
@@ 2.1 The line as an effective description
and infinite potential walls at 0 and $L$. Say we are mainly interested in a state supported near $x=L/2$, dying off very quickly relative to $L$, such as (schematically)
@@
[…]
@@ 2.2 Codifying the relationship with $V$ and $R^{*}$
[…]
@@
[…]
@@
[…]
$\braket{\psi|O|\psi}\approx\braket{\psi|V_{\varepsilon}^{\dagger}R^{*}(O)V_{\varepsilon}|\psi}$ (2.14)
@@ 2.3 Lessons for JT
Recall that one idea, from [Iliesiu:2024cnh], is
@@
[…]
@@
Note a crucial difference between JT and this line/interval example. In JT, we are starting by knowing $V$ and would like to deduce $R^{*}$. This is trickier than the converse, because $V$ can be straightforwardly determined from $R$ via (2.10) and (2.11). We’$l$ have to figure out the appropriate $R^{*}$ by different means.